### v1

# Latent Topic Synthesis: Leveraging LLMs for Electoral Ad Analysis

Abstract
Social media platforms play a pivotal role in shaping political discourse, but analyzing their vast and rapidly evolving content remains a major challenge. We introduce an end-to-end framework for automatically generating an interpretable topic taxonomy from an unlabeled corpus. By combining unsupervised clustering with prompt-based labeling, our method leverages large language models (LLMs) to iteratively construct a taxonomy without requiring seed sets or domain expertise. We apply this framework to a large corpus of Meta (previously known as Facebook) political ads from the month ahead of the 2024 U.S. Presidential election. Our approach uncovers latent discourse structures, synthesizes semantically rich topic labels, and annotates topics with moral framing dimensions. We show quantitative and qualitative analyses to demonstrate the effectiveness of our framework. Our findings reveal that voting and immigration ads dominate overall spending and impressions, while abortion and election-integrity achieve disproportionate reach. Funding patterns are equally polarized: economic appeals are driven mainly by conservative PACs, abortion messaging splits between pro- and anti-rights coalitions, and crime-and-justice campaigns are fragmented across local committees. The framing of these appeals also diverges—abortion ads emphasize liberty/oppression rhetoric, while economic messaging blends care/harm, fairness/cheating, and liberty/oppression narratives. Topic salience further reveals strong correlations between moral foundations and issues, such as fairness/cheating with crime/justice and loyalty/betrayal with immigration. Demographic targeting also emerges: younger Floridians are shown affordable-housing ads while older groups receive abortion-focused appeals; Montana males are disproportionately targeted with environmental and liberty-oriented messaging, whereas Virginia males are mobilized around crime/justice and voting. This work supports scalable, interpretable analysis of political messaging on social media, enabling researchers, policymakers, and the public to better understand emerging narratives, polarization dynamics, and the moral underpinnings of digital political communication.
Introduction
Modern political campaigns have been fundamentally transformed by the rise of social media platforms. Platforms such as Meta have become critical battlegrounds for political entities, enabling them to target specific voter segments with tailored messages. Chu et al. (2024) has shown that voters respond more favorably to ads on issues they personally prioritize, irrespective of partisanship. Understanding these targeting strategies is thus essential for both researchers and policymakers. As the volume and velocity of such dynamic content continues to grow, there is an urgent need for scalable tools that can systematically analyze these messages.
Figure 1: Framework overview. All ads are initially embedded and clustered. (a) Each cluster is sequentially processed by the LLM to extend the topic taxonomy. (b) The generated topics are used to label the ads in each cluster.
Manual annotation of large-scale text corpora is expensive and time-consuming, prompting widespread use of topic modeling techniques such as Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) and non-negative matrix factorization (NMF) (Lee and Seung 1999) for unsupervised topic discovery. Although effective in identifying topics, these methods produce topic representations that are not readily interpretable and often require human labeling. To automate this step, previous works have explored probabilistic labeling (Magatti et al. 2009), summarization-based methods (Wan and Wang 2016; Cano Basave, He, and Xu 2014), and the sequence-to-sequence model (Alokaili, Aletras, and Stevenson 2020). Recently, hierarchical topic modeling frameworks such as BERTopic (Grootendorst 2022) have gained wide adoption. BERTopic builds on contextualized representations and density-based clustering. Building on this approach, we propose a multi-step topic generation framework that leverages large language models (LLMs) (Brown et al. 2020) to produce semantically rich and coherent labels.
Recent studies have also explored the use of LLMs to identify nuanced topics and uncover latent themes and arguments. LLMs-in-the-loop approach has been used for understanding fine-grained topics (Islam and Goldwasser 2025c, b). Islam and Goldwasser (2025b) employed a seed set of initial themes to guide their framework, while Islam and Goldwasser (2025c) assumed a predefined set of themes and then employ their framework to uncover arguments. In contrast, our method operates without any initial seed set, instead iteratively constructing a topic taxonomy from scratch. This enables greater flexibility and adaptability in capturing latent discourse structures within large, dynamic social media corpora.
In this paper, we propose a novel two-pass iterative topic generation framework (Figure 1) that combines the strengths of unsupervised machine learning with the interpretive capabilities of LLMs. Our approach applies unsupervised clustering techniques to identify latent topical structures within a document corpus, followed by an iterative process where LLMs evaluate existing topics and generate new ones as needed. This method allows for the discovery of semantically rich and coherent topics without relying on predefined labels or human intervention. The output of this process is a set of coherent and interpretable document clusters, which can be used for downstream tasks such as supervised classification.
We demonstrate the effectiveness of this approach through a comprehensive case study [^1] Code and dataset will be released upon publication. of Meta political advertisements, taken one month before the 2024 US Presidential Election. We uncover a diverse political issue taxonomy that captures the topical landscape of the electoral campaign. We further extend our analysis by examining the moral framings underlying these ads, under the lens of Moral Foundation Theory (Haidt and Graham 2007; Haidt and Joseph 2004). By analyzing how different political actors segment and appeal to voter demographics, our work provides valuable insights into contemporary campaign tactics while offering a methodological contribution that bridges traditional topic modeling with the interpretive capabilities of LLMs. Our contributions include:
  1. 1.
    A flexible framework for interpretable topic modeling that requires no human intervention.
  2. 2.
    A dynamic taxonomy of political issues and moral foundations, generated from unlabeled data with no seed set.
  3. 3.
    Systematic analysis of targeting strategies in political social media advertising.
  4. 4.
    Curate and release the 2024 US presidential election dataset to support future research.
Related Work
Social media campaigns are a powerful force in shaping public discourse (Islam 2025b, a; Islam and Goldwasser 2024; Islam, Roy, and Goldwasser 2023; Islam, Zhang, and Goldwasser 2023; Islam and Goldwasser 2022; Capozzi et al. 2021; Goldberg et al. 2021), influencing elections (Ribeiro et al. 2019; Silva et al. 2020; Fulgoni, Lipsman, and Davidsen 2016), and mobilizing voters (Aggarwal et al. 2023; Teresi and Michelson 2015; Hersh 2015a). Research on political microtargeting (Tappin et al. 2023; Zuiderveen Borgesius et al. 2018; Hersh 2015b) has been a key focus, with studies examining how targeted ads can influence voter preferences and behavior (Hirsch et al. 2024).
Researchers have explored the role of targeted messaging in elections and the strategic dissemination of campaign content across platforms (Hackenburg and Margetts 2024; Islam, Roy, and Goldwasser 2023). These studies underscore the growing need for scalable methods to analyze social media content at scale and in context, particularly during major political events such as presidential campaigns.
However, the sheer volume of social media data presents a significant challenge for researchers and practitioners. Traditional manual coding methods are often impractical for large datasets, prompting widespread use of topic modeling techniques such as LDA across various domains, including social media (Chen et al. 2019; Ozer, Kim, and Davulcu 2016) and political discourse (Lahoti, Garimella, and Gionis 2018; Mathaisel and Comm 2021). These methods have been instrumental in uncovering topics in text data, but they often struggle to produce coherent and interpretable topics (Chang et al. 2009; Mei, Shen, and Zhai 2007).
To address these issues, modern embedding-based topic models like BERTopic (Grootendorst 2022) utilize contextualized representations and density-based clustering, resulting in more coherent and interpretable topics. Our framework builds on these advances by following the density-based clustering approach of BERTopic, but relying on LLM inference to generate an interpretable topic taxonomy.
Prompt-based methods such as TopicGPT (Pham et al. 2023) have been shown to outperform traditional topic modeling techniques in terms of coherence and interpretability. These methods leverage LLMs to synthesize topic taxonomies and assign labels directly, allowing for more nuanced and context-sensitive analysis. Notably, Gilardi, Alizadeh, and Kubli (2023) showed that LLMs match or exceed the performance of expert human annotators in labeling tasks.
More recently, LLMs-in-the-loop approaches have been used for understanding fine-grained topics (Islam and Goldwasser 2025c, b). In their earlier work, Islam and Goldwasser (2025b) guided their framework using a seed set of initial themes. Building on this, Islam and Goldwasser (2025c) assumed a predefined set of themes and focused on uncovering underlying arguments. In contrast, our method operates without any initial seed set.
Another growing area of interest is the moral framing of political messages, often analyzed through the lens of Moral Foundations Theory (MFT) (Haidt and Graham 2007; Haidt and Joseph 2004). Earlier work relied on manually annotated datasets to train classifiers that detect moral appeals in text (Pacheco, Islam et al. 2022; Roy, Pacheco, and Goldwasser 2021). Recently, Islam and Goldwasser (2025a) has leveraged LLMs to generate moral labels and explanations through inference, allowing for more scalable and nuanced analysis. We extends this line of research by employing LLMs to synthesize arguments at the cluster level and assign corresponding moral labels, resulting in a more contextualized and semantically grounded interpretation of political discourse.
Our work builds on these advances by introducing an unsupervised, LLM-guided framework for analyzing political ads. Without relying on pre-defined themes or seed sets, we ensure that the generated topics are coherent and interpretable. This enables scalable, and context-sensitive analysis of political campaigns on social media: an increasingly vital task in the era of digital microtargeting.
Methodology
We propose a framework for the automated annotation of large datasets of unstructured text data, for example, social media political advertisements. This framework consists of three key components: (1) embedding-based clustering, (2) large language model (LLM) topic synthesis, (3) LLM-based annotation.
Issue #Clusters Example Ad
economy 15 Molly Buck’s agenda is failing Iowa families. We can’t afford Molly Buck in the State House.
voting rights 13 Make Your Vote Count. RE-register to Vote in the District of Your Second Home.
crime/justice 12 Orange County Firefighters Trust Dave Min To Keep Our Communities Safe.
personal freedom 4 Tim Sheehy will always fight for Montana Values in the Senate!
voting 4 Vote Rebecca for State Rep by Nov 5th!
education 4 Vote for students. Vote for teachers. Support our public schools.
affordable housing 4 Catherine has supported more multi-family housing so we can afford homes in the communities we grew up in. Vote Stefani!
environmental protection 4 We can count on Lucy Rehm to be a responsible leader and protect our clean air and water.
immigration 3 Yadira Caraveo supported efforts to defund ICE and the border patrol. Now Colorado is being overwhelmed with illegal immigration.
election integrity 3 Question 3 will change our state’s Constitution and weaken our democracy by throwing away thousands of ballots over minor errors. Vote NO on Question 3.
abortion 2 Defend reproductive freedom - VOTE on Tuesday, November 5th!
healthcare access 1 Lucy Rehm cares about improving conditions in your local hospital. That’s why Minnesota Nurses endorsed Lucy Rehm!
property taxes 1 Caleb will work to stop reckless spending and oppose tax and fee increases.
other 2 Get the facts on the Gloucester Township sewer sale, visit VoteYesGloucesterTownship.com!
Table 1: Topics identified by the LLM in the political ads dataset. Clusters column gives the number of unique clusters assigned to each topic by the LLM.
Density-based Clustering
The input documents are first embedded into a high-dimensional space using a pre-trained sentence embedding model, such as Sentence-BERT (Reimers and Gurevych 2019). To avoid the curse of dimensionality, we use UMAP (McInnes, Healy, and Melville 2020) to reduce the dimensionality of the embeddings to a more manageable size. UMAP is a non-linear dimensionality reduction technique that preserves the local structure of the data, making it suitable for high-dimensional data such as text embeddings.
We then use HDBSCAN (McInnes, Healy, and Astels 2017), a hierarchical density-based clustering algorithm, to group similar text data points together and identify topical structures in the data. This is particularly useful for unstructured text data, where the distribution of data points may not follow a standard distribution and noise may be present. For each cluster, we extract the five items with the highest membership probabilities, which we refer to as the cluster representatives.
Iterative Topic Synthesis
To generate a domain-specific label taxonomy, we use an LLM to synthesize labels from the clusters. The initial label list is set empty, or optionally initialized with a small seed set to guide the generation process. This list will be iteratively expanded by the LLM. If the initial list is empty, the LLM is prompted to generate a label for the first cluster.
For every other cluster, we prompt the LLM whether any of the existing labels are accurate annotations for the cluster representatives. We use constrained decoding (Beurer-Kellner, Fischer, and Vechev 2024) to force the LLM to only answer with “yes” or “no”. If the LLM answers “no”, we prompt it to generate a new label for the cluster. We then add this label to the set of labels and repeat the process for all clusters. This iterative process continues until all clusters have been processed.
Cluster Representative Labeling
The next step is to use the generated label taxonomy to annotate the cluster representatives. We prompt the LLM to generate a label for each set of cluster representatives. We once again utilize constrained decoding to limit the LLMs output to only one of the possible labels.
Supervised Classification
A downstream supervised classification task can optionally be performed using the cluster representatives and their assigned labels. This is particularly useful as clusters can be noisy, and HDBSCAN does not assign each (or even most) data points to a cluster.
For classification, we found SetFit (Tunstall et al. 2022) to be particularly effective for text classification tasks with limited labeled data. SetFit uses a two-step process: first, it fine-tunes a pre-trained sentence transformer model on the labeled data points using a contrastive loss function. Second, it trains a classification head to map the fine-tuned embeddings to the label space. This model is then used to efficiently annotate the remaining unlabeled documents.
Case Study
To demonstrate the effectiveness of our approach, we apply it to a large corpus of social media political advertisements from the 2024 US presidential elections. The goal is to uncover patterns the advertisers employ to target specific demographics with certain issues and moral framings.
Care/Harm: : It suggests that someone other than the speaker is worthy of compassion or is experiencing harm, grounded in the values of kindness, tenderness, and care.
Fairness/Cheating: Emphasizes justice, personal rights, and independence; involves comparing with other groups. Advocates for equal opportunity and resists those who benefit without contributing (“Free Riders”).
Loyalty/Betrayal: Based on the values of loyalty to one’s country and willingness to sacrifice for the group. Activated by a sense of unity and collective responsibility—“one for all, and all for one”.
Authority/Subversion: Centers on showing respect (or resistance) toward established authority and following long-standing traditions. It includes maintaining social order and fulfilling the duties tied to hierarchical roles, such as obedience, respect, and role-based responsibilities.
Sanctity/Degradation: Beyond religion, this value highlights respect for human dignity and aversion to moral or physical corruption, promoting purity, self-control, and the belief that the body is sacred and vulnerable to defilement.
Liberty/Oppression: Captures the feelings of reactance and resentment people experience when their freedom is restricted, often leading to collective disdain for authoritarian figures and motivating unity and resistance against oppression.
Table 2: Six basic moral foundations (Haidt and Graham 2007; Haidt and Joseph 2004).
Dataset
We collected a dataset of political advertisements running on Facebook and Instagram in the USA from October 2024. These ads were collected from the Meta Ad Library and contains $8047$ unique ads. Crucially, these ads were chosen on a day only a month away from the 2024 Presidential election, with the ad’s content covering election-related topics at both the local and federal levels. Over $99\%$ of the ads are in English, with a small fraction in Spanish, mostly in local Florida elections.
For each ad, Meta provides a unique ad ID, the ad title, ad body, and URL, ad creation time and the time span of the campaign, the Meta page authoring the ad, funding entity, and the cost of the ad (given as a range). The API also provides information on the users who have seen the ad (called ‘impressions’): the total number of impressions (given as a range and we take the average of the end points of the range), distribution over impressions broken down by gender (male, female, unknown), age ($7$ groups), and location down to the state level.
Framework Application
We apply the proposed framework to the ad texts to identify a taxonomy of political issues discussed in the ads. We embedded the ad body texts using the all-MiniLM-L6-v2 model, and removed all ads with embedding vectors too similar to another ad (cosine similarity >0.95>0.95). We then clustered these embeddings with HDBSCAN using a minimum cluster size of $15$. This resulted in $4510$ of the ads being assigned to $72$ clusters. The prompts can be found in the Appendix A.
We used a locally hosted Llama-3.2-3B model (Grattafiori et al. 2024) to generate the cluster labels. Experiments were conducted on a local machine using an 11th Gen Intel Core i7-11390H CPU @ 3.40GHz, without GPU acceleration. We began the topic synthesis process without any seed set, which the model expanded to $14$ total topics. For annoation, the model was additionally given the option to assign a cluster the label ‘other’ if none of the topics suited its context. Of this list, the LLM never assigned the topic ‘border security’ to any of the clusters. See Table 1 for the final list of topics and the number of clusters assigned to each topic.
Moral Foundations
In addition to topic assignment, we classify the moral foundation of each ad. Moral Foundation Theory (MFT) suggests a theoretical framework for analyzing six moral values (i.e., foundations, each with a positive and a negative polarity) central to human moral sentiment (Table 2). MFT states that political attitudes are shaped by six core moral dimensions: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, sanctity/degradation, and liberty/oppression.
To identify the moral foundation of each clusters, we first prompt the LLM to summarize the primary talking point of the cluster representatives with regards to the annotated label. The LLM is then provided with definitions of the six moral foundations, and the extracted argument is classified using constrained decoding to determine the most strongly aligned foundation. Prompts are provided in the Appendix.
Supervised Classification
To annotate the unlabeled ads, we use the labeled cluster representatives as the training set for a supervised classification task. We used a variety of classifiers, including Logistic Regression, XGBoost (Chen and Guestrin 2016), RoBERTa (Liu et al. 2019), and SetFit (Tunstall et al. 2022).
For the labeled ads, we took all ads which HDBSCAN assigned a high membership probability to (greater than $0.98$), which resulted in a total of $2337$ ads, of which the four majority classes ‘voting rights’ (26%), ‘crime/justice’ (19%), ‘education’ (17%), and ‘abortion’ (11%) made up 3 quarters (73%) of the labeled ads. The remaining classes were much smaller, with the smallest class being ‘healthcare access’ and ‘property taxes’ (0.64% each).
We randomly selected $200$ ads from the unassigned clusters, and annotated them manually. We used these annotations as the ground truth labels for the classification task.
For each model, we used the all-MiniLM-L6-v2 model to embed the ad texts and employed GridSearchCV to optimize the hyperparameters. We used $5$-fold cross-validation to evaluate the models, and selected the best performing model based on the macro F1 score. The results of this evaluation are shown in Table 3.
Model Macro F1 Accuracy
Logistic Regression 0.37 0.55
XGBoost 0.31 0.53
RoBERTa 0.32 0.53
SetFit 0.36 0.6
Table 3: Classification performance across annotated ads for topic assignment (15 unbalanced classes).
Demographic Targeting Analysis
Using the ad impressions data, we can analyze the targeting data of the ads. The analysis focuses on the use of social media platforms, particularly Meta and Instagram, to disseminate targeted political advertisements. We examine the strategies employed by the advertisers to reach specific demographic groups.
Meta allows advertisers to define their audiences on the basis of location (state-level), age and gender. We utilized this data, together with the annotations, to analyze the microtargeting strategies of the political entities.
To understand how messages are tailored to different demographics, we looked at the positive pointwise mutual information (PPMI) between the words in the advertisements and the demographic groups. The PPMI is a measure of association between two events, in this case, the words in the advertisements and the demographic groups. A higher PPMI value indicates a stronger association between a word and a demographic group. We compared both the topics discussed in the advertisements as well as the moral foundations that were used to appeal to the different demographic groups. This annotated corpus enabled detailed analysis of issue emphasis and demographic targeting, which we report in the Results and Analysis Section.
Results and Analysis
We evaluate the performance of our approach on the social media advertisement dataset, evaluating the performance of both annotation and classification tasks. Additionally, we discuss the results of our case study, showcasing differences in moral framing and issue distribution across different demographics.
Model Average Score Best Label
BERTopic 1.2 3
Our Method 2.8 12
Table 4: Annotation results. The average score is out of 5, and the best label is the number of times that model’s label was selected as the most fitting (among all annotators).
Figure 2: Average spend and impressions reveal which topics dominate the ad landscape.
Figure 3: Top five funding entities by total spend across economy, abortion, and crime/justice topics. Economic ads are dominated by conservative groups, abortion spending reflects polarized investments from both anti-abortion and pro-abortion rights organizations, and crime/justice advertising is driven by a mix of PACs and local campaigns.
Baseline Comparison
We evaluated the topic labels, comparing the performance of our proposed framework with the baseline BERTopic  (Grootendorst 2022). For both models, the same clusters were used, and both models used the same LLM to generate the topic labels. For BERTopic, the LLM generated the labels based on both the representative documents and the top-10 keywords.
Two annotators were each given 10 randomly selected clustered ads. The annotators were volunteer graduate students who spoke fluent English with backgrounds in Computational Social Science and Computer Science. For each ad, they were given both the label from BERTopic and our proposed framework (the order of the labels was randomized). The annotators were asked to select the label that best represented the ad, and to provide a score from 1-5 for each label, where 5 would be a perfect fit, and 1 would describe a totally irrelevant label. The results of this comparison are shown in Table 4.
These results show that our proposed framework outperforms BERTopic in terms of both average score and the number of times the label was selected as the best label. A Cohen’s Kappa (Cohen 1960) score of $0.66$ was achieved between the two annotators, indicating a moderate level of agreement. For $5$ out of the $20$ rounds, the annotators gave the same score for both models (in all cases a score of 1). The average score of $2.8$ for our proposed framework indicates that the labels generated are more relevant and accurate compared to the average score of $1.2$ for BERTopic.
Spending and Reach by Topic
We begin with a descriptive analysis of ad spending and exposure across topics. Figure 2 shows the average spend (left panel) and average impressions (right panel) per topic. Overall, voting and immigration dominate both spending and impressions, while issues such as healthcare access and property taxes receive relatively little paid attention. Interestingly, abortion and election integrity attract intermediate spending yet generate disproportionally high impressions, suggesting higher engagement or lower cost per impression in these issue domains.
Figure 4: Topic-specific moral narratives: multiple frames shape economic ads, Liberty/Oppression dominates abortion related ads.
Top Funders by Issue Domain
We examine the concentration of spending among funding entities. Figure 3 presents the top $5$ funding entities by total spend for one of the three central issues [^2] https://www.pewresearch.org/politics/2024/09/09/issues-and-the-2024-election/ in the 2024 election: economy, abortion, and crime/justice.
On the economy, conservative-aligned groups dominate the funding landscape. The Keystone [^3] https://www.keystonepac.org/ and the Congressional Leadership Fund [^4] https://congressionalleadershipfund.org/ lead spending, followed by organizations such as Americans for Prosperity Action [^5] https://americansforprosperity.org/. This pattern reflects Republican priorities around fiscal and economic messaging, emphasizing taxation, regulation, and inflation.
For abortion, spending is highly polarized. Florida Voters Against Extremism [^6] https://dos.elections.myflorida.com/committees/ComDetail.asp?account=84315 is the largest spender, anchoring anti-abortion rights mobilization, particularly around state-level ballot initiatives. In contrast, progressive-aligned group such as Montanans Securing Reproductive Rights  [^7] https://www.influencewatch.org/organization/montanans-securing-reproductive-rights/ invests heavily in support of abortion rights. This split underscores abortion as a wedge issue, with both sides committing substantial but asymmetrically distributed resources.
In the case of crime and justice, the funding environment is distinct. Judicial Fairness PAC [^8] https://judicialfairnesspac.org/ (with messages focused on judicial appointments and criminal justice reform) and Smart and Safe Florida [^9] https://smartandsafeflorida.com/ (with messages focused on public safety) together account for a large share of spending. Other contributors, such as Stop Houston Murders PAC [^10] https://stophoustonmurders.com/ and Iowans for Zach Nunn [^11] https://www.fec.gov/data/committee/C00784389/, highlight localized campaign investments. Notably, this issue domain demonstrates geographically fragmented but financially intensive spending, reflecting its resonance in state and local politics as well as in national-level partisan framing.
Overall, Figure 3 shows that while the economy and abortion are dominated by major national organizations with strong partisan alignment, crime/justice advertising is more decentralized, fueled by a mix of PACs and candidate-aligned committees targeting specific jurisdictions.
Moral Foundations in Funded Messaging
We link moral framing to financial investment. Figure 4 compares the top $5$ funding entities in two salient topics — economy and abortion — broken down by their dominant moral foundation appeals. On the economy, funding entities split their resources between Care/Harm frames (e.g., worker hardship) and Fairness/Cheating or Liberty/Oppression frames (e.g., taxation, regulation). On abortion, almost all spending is channeled through a Liberty/Oppression frame, consistent with rhetoric around government interference in reproductive rights. This illustrates that moral narratives are not evenly distributed across topics: economic appeals are morally plural, while abortion appeals are morally singular.
Topic Salience and Moral Framing
We present some selected results from our case study. Firstly, we analyze the correlation between moral foundations and the latent topics, taking LDA topics as a baseline. To compute the correlation, we first one-hot encoded each ad twice: once by topic and once by moral foundation. We then computed the pairwise correlation using the Pearson correlation between these one-hot encodings. The results are shown in Figure 5. In Figure 5(a), we can see very strong correlations between certain moral foundations and topics. For example, the Fairness/Cheating moral axis is strongly correlated with the crime/justice topic, while the Loyalty/Betrayal moral axis is strongly correlated with the immigration topic. This suggests that our proposed framework is able to capture the moral framing of the advertisements. In contrast, the LDA topics in Figure 5(b) show much weaker correlations with the moral foundations.
(a) Proposed Framework Topics
(b) LDA Topics
Figure 5: Correlations between moral foundations and topics. Color intensity indicates the strength of the correlation.
(a) Topics: Females in Alabama vs Washington
(b) Topic: All Age groups in Florida
(c) Topic: Males in Montana vs Virginia
(d) Moral Framing across States
Figure 6: Heatmaps showing the PPMI analysis of moral framings and political issues across different demographics. The color intensity indicates the strength of the correlation.
We also analyzed the distribution of moral framings and political issues across different demographics to uncover political microtargeting strategies. We used the positive point-wise mutual information (PPMI) to detect topics that appear disproportionately in ads shown to specific demographics, potentially indicating targeted messaging. We visualize the results of the PPMI analysis in Figure 6. The heatmaps show the distribution of moral framings and political issues across different demographics. Some interesting observations include the differences in ads targeting different states, or the shift in topics among different age groups in Florida (Figure 6(b)). For example, young people in Florida are more likely to be targeted with ads related to affordable housing, while older people are more likely to be targeted with ads related to abortion. This shows generational targeting strategies, where economic concerns are emphasized for younger groups, while older groups are mobilized on moralized issues like abortion.
On the other hand, Figure 6(c) shows that environmental protection and personal freedom are heavily overrepresented among male audiences in Montana, while abortion also appears moderately elevated. By contrast, male audiences in Virginia are disproportionately exposed to crime/justice and voting-related advertising. These contrasts highlight how campaign strategies selectively target localized demographic niches: progressive-leaning issue frames resonate more strongly in Montana male audiences, whereas mobilization around law-and-order and electoral process is more pronounced among Virginia males.
Conclusion
We produce a modular framework for large-scale annotation of unstructured text documents. While our framework is generalizable, the key contribution here is enabling large-scale, interpretable analysis of electoral ads. Applied to a corpus of eight thousand political advertisements, our method enabled efficient labeling with minimal human effort and low computational cost. We demonstrate that our approach is effective for both supervised and unsupervised learning tasks, and the analysis of the resulting annotations uncovered meaningful patterns in issue framing across demographic groups. In addition, the release of a new dataset will provide a valuable resource for the research community. We believe our framework can be applied to a variety of text classification tasks, discovering latent structures in large corpora, and enable cheap and efficient automated annotation of unstructured documents. Our findings demonstrate that combining clustering with LLM-based labeling yields scalable, interpretable analysis of political ads, surfacing both substantive issue agendas and the moral frames used to mobilize audiences. These insights show how social media advertising both reflects and amplifies polarization, providing researchers and policymakers with tools to monitor emerging narratives and targeting strategies.
Limitations
While our framework reduces manual annotation effort and generalizes across domains, several limitations remain. Firstly, we purposefully choose a weaker LLM to show that our framework is not limited to computational resources or budget constraints. However, the performance of our framework is likely to improve with more powerful LLMs. Secondly, while we demonstrate the effectiveness of our framework on a political advertisement corpus, it may not generalize to all domains.
Our framework relies on LLMs, which may introduce biases or hallucinate misleading topic labels. Additionally, automatic interpretation of political content may risk oversimplification or misrepresentation of nuanced discourse.
Future work should explore the applicability of our framework across different domains and tasks. Clustering performance can be sensitive to embedding quality and density-based parameters, and we recommend further investigation into the optimal parameters for different datasets.
Ethics Statement
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of the results. The author’s personal views are not represented in any results we report, as it is solely outcomes derived from machine learning and/or AI models. The data collected in this work was made publicly available by the Meta Ad Library API. The data do not contain personally identifiable information and report engagement patterns at an aggregate level.
Appendix A Prompts
Below are the prompts used in the various steps of the framework. Constrained decoding was implemented with the guidance framework [^12] https://github.com/guidance-ai/guidance.
System Prompt
You are a political analyst. You are given the following set of political ads to analyze. You will assign the ads a topic that best summarizes the key issue discussed in these ads.
Topic Synthesis
Binary Output
Is there a topic in the following list that well describes the key issue discussed in these ads?
Topics:
{{ topics }}
Can the ads can be summarized by one of the previous topics: Answer with ”yes” or ”no”.
Generate New Topic (User)
In three words or less, describe the issue that ads discuss, summarizing the topic. Examples: ”Abortion”, ”Climate Change”, etc.
Output constriained to [”yes”, ”no”].
Generate New Topic (Assistant)
A better topic for these ads is: ”
The LLM would finish the above, with the quote (”) as a stop token.
Annotation
System Prompt
You are a political analyst. You are given the following set of political ads to analyze: {{ ads }}
Select Topic (User)
Summarize the main talking point of the ads. Do so in a single sentence.
Output constrained to set of topics.
Summarize (User)
Summarize the main talking point of the ads. Do so in a single sentence.
Select Moral Foundation (User)
Which of Which of the following moral foundations best describes the arguments used in the ads about {{ topic }}?’
{{ moral_foundations_with_definitions }}