### v2

# Iterative Topic Taxonomy Induction with LLMs: A Case Study of Electoral Advertising

Abstract
Social media platforms play a pivotal role in shaping political discourse, but analyzing their vast and rapidly evolving content remains a major challenge. We introduce an end-to-end framework for automatically inducing an interpretable topic taxonomy from unlabeled text corpora. By combining unsupervised clustering with prompt-based inference, our method leverages large language models (LLMs) to iteratively construct a taxonomy without requiring seed sets (predefined labels) or domain expertise. We validate the framework through a study of political advertising ahead of the 2024 U.S. presidential election. The induced taxonomy yields semantically rich topic labels and supports downstream analyses, including moral framing, in this setting. Results suggest that structured, iterative labeling yields more consistent and interpretable topic labels than existing approaches under human evaluation, and is practical for analyzing large-scale political advertising data.
Iterative Topic Taxonomy Induction with LLMs: A Case Study of Electoral Advertising
Alexander Brady1thanks: These authors contributed equally. 1Department of Computer Science, ETH Zürich, Zurich, Switzerland bradya@ethz.ch          Tunazzina Islam211footnotemark: 1 2Department of Computer Science, Purdue University, West Lafayette, IN 47907, USA islam32@purdue.edu
Figure 1: Framework overview. All ads are initially embedded and clustered. (a) Each cluster is sequentially processed by the LLM to extend the topic taxonomy. (b) The generated topics are used to label the ads in each cluster.
1 Introduction
Modern political campaigns have been fundamentally transformed by the rise of social media platforms. Platforms such as Meta (previously known as Facebook) have become critical battlegrounds for political entities, enabling them to target specific voter segments with tailored messages. Chu et al. (2024) have shown that voters respond more favorably to ads on issues they personally prioritize, irrespective of partisanship. Understanding these targeting strategies is thus essential for both researchers and policymakers. As the volume and velocity of such dynamic content continues to grow, there is an urgent need for scalable tools that can systematically analyze such messages at scale and with minimal human supervision.
Manual annotation of large-scale text corpora is expensive and time-consuming, prompting widespread use of topic modeling techniques such as Latent Dirichlet Allocation (LDA) Blei et al. (2003) and non-negative matrix factorization (NMF) Lee and Seung (1999) for unsupervised topic discovery. Although effective in identifying topics, these methods produce topic representations that are not readily interpretable and often require human labeling. To automate this step, previous works have explored probabilistic labeling Magatti et al. (2009), summarization-based methods Wan and Wang (2016); Cano Basave et al. (2014), and the sequence-to-sequence model Alokaili et al. (2020). Recently, hierarchical topic modeling frameworks such as BERTopic Grootendorst (2022) have gained wide adoption. BERTopic builds on contextualized representations and density-based clustering. Building on this approach, we propose a multi-step topic generation framework that leverages large language models (LLMs) Brown et al. (2020) to produce semantically rich and coherent labels.
Recent studies have also explored the use of LLMs to identify nuanced topics and uncover latent themes and arguments. LLMs-in-the-loop approach has been used for understanding fine-grained topics Islam and Goldwasser (2025d, b). Islam and Goldwasser (2025b) employed a seed set of initial themes to guide their framework, while Islam and Goldwasser (2025d) assumed a predefined set of themes and then employ their framework to uncover arguments. In contrast, our method operates without any initial seed set, instead iteratively constructing a topic taxonomy from scratch. This enables greater flexibility and adaptability when inducing topic structure in large, dynamic social media corpora.
In this paper, we propose a novel two-pass iterative topic generation framework (Figure 1) that combines the strengths of unsupervised machine learning with the interpretive capabilities of LLMs. Our approach applies unsupervised clustering techniques to identify latent topical structures within a document corpus, followed by an iterative process where LLMs evaluate existing topics and generate new ones as needed. This method allows for the discovery of semantically rich and coherent topics without relying on predefined labels or direct human intervention. The output of this process is a set of coherent and interpretable document clusters, which can be used for downstream tasks such as supervised classification.
We demonstrate the effectiveness of this approach through a comprehensive case study of Meta political advertisements, taken one month before the 2024 US Presidential Election [^1] Code and dataset are publicly available here: https://github.com/alexander-brady/llm-topic-synthesis. We uncover a diverse political issue taxonomy that captures the topical landscape of the electoral campaign. We further extend our analysis by examining the moral framings underlying these ads, under the lens of Moral Foundation Theory Haidt and Graham (2007); Haidt and Joseph (2004). By analyzing how different political actors segment and appeal to voter demographics, the case study illustrates the types of substantive insights enabled by the proposed framework. Our contributions include:
  1. 1.
    A general, seed-free, LLM-guided framework for interpretable topic taxonomy induction from unlabeled corpora without human intervention.
  2. 2.
    An iterative topic synthesis mechanism that yields more consistent and better-aligned topic labels than classical topic models and single-shot LLM labeling under human evaluation.
  3. 3.
    A real world political advertising case study used to validate the framework and illustrate its analytical affordances.
  4. 4.
    Release of a curated dataset of 2024 U.S. presidential election ads to support future research.
2 Related Work
Social media campaigns are a powerful force in shaping public discourse (Islam, 2025b, a; Islam and Goldwasser, 2025c; Islam et al., 2023a, b; Islam and Goldwasser, 2022; Capozzi et al., 2021; Goldberg et al., 2021), influencing elections (Ribeiro et al., 2019; Silva et al., 2020; Fulgoni et al., 2016), and mobilizing voters (Aggarwal et al., 2023; Teresi and Michelson, 2015; Hersh, 2015a). Research on political microtargeting Tappin et al. (2023); Zuiderveen Borgesius et al. (2018); Hersh (2015b) has been a key focus, with studies examining how targeted ads can influence voter preferences and behavior (Hirsch et al., 2024).
Researchers have explored the role of targeted messaging in elections and the strategic dissemination of campaign content across platforms Hackenburg and Margetts (2024); Islam et al. (2023a). These studies underscore the growing need for scalable methods to analyze social media content at scale and in context, particularly during major political events such as presidential campaigns.
However, the sheer volume of social media data presents a significant challenge for researchers and practitioners. Traditional manual coding methods are often impractical for large datasets, prompting widespread use of topic modeling techniques across various domains, including social media Chen et al. (2019); Ozer et al. (2016) and political discourse Lahoti et al. (2018); Mathaisel and Comm (2021). These methods have been instrumental in uncovering topics in text data, but they often struggle to produce coherent and interpretable topics Chang et al. (2009); Mei et al. (2007).
To address these issues, modern embedding-based topic models like BERTopic Grootendorst (2022) utilize contextualized representations and density-based clustering, resulting in more coherent and interpretable topics. Our framework builds on these advances by following the density-based clustering approach of BERTopic, but relying on LLM inference to generate an interpretable topic taxonomy.
Prompt-based methods such as TopicGPT Pham et al. (2023) have been shown to outperform traditional topic modeling techniques in terms of coherence and interpretability. These methods leverage LLMs to synthesize topic taxonomies and assign labels directly, allowing for more nuanced and context-sensitive analysis. Notably, Gilardi et al. (2023) showed that LLMs match or exceed the performance of expert human annotators in labeling tasks.
More recently, LLMs-in-the-loop approaches have been used for understanding fine-grained topics Islam and Goldwasser (2025d, b). In their earlier work, Islam and Goldwasser (2025b) guided their framework using a seed set of initial themes. Building on this, Islam and Goldwasser (2025d) assumed a predefined set of themes and focused on uncovering underlying arguments. Lam et al. (2024) developed a concept induction algorithm using LLM with human-guided abstraction called LLooM, which has a seed operator for accepting user-provided seed term/set. In contrast, our method operates without any initial seed set.
Another growing area of interest is the moral framing of political messages, often analyzed through the lens of Moral Foundations Theory (MFT) Haidt and Graham (2007); Haidt and Joseph (2004). Earlier work relied on manually annotated datasets to train classifiers that detect moral appeals in text Pacheco et al. (2022); Roy et al. (2021). Recently, Islam and Goldwasser (2025a) has leveraged LLMs to generate moral labels and explanations through inference, allowing for more scalable and nuanced analysis. We extend this line of research by employing LLMs to synthesize arguments at the cluster level and assign corresponding moral labels, resulting in a more contextualized and semantically grounded interpretation of political discourse.
Our work builds on these advances by introducing an unsupervised, LLM-guided framework for analyzing political ads. Without relying on pre-defined themes or seed sets, we ensure that the generated topics are coherent and interpretable. This enables scalable, and context-sensitive analysis of political campaigns on social media: an increasingly vital task in the era of digital microtargeting.
3 Methodology
We propose a framework for the automated annotation of large datasets of unstructured text data, for example, social media political advertisements. This framework consists of three key components: (1) embedding-based clustering, (2) large language model (LLM) topic synthesis, (3) LLM-based annotation.
3.1 Density-based Clustering
The input documents are first embedded into a high-dimensional space using a pre-trained sentence embedding model, such as Sentence-BERT (Reimers and Gurevych, 2019). To avoid the curse of dimensionality, we use UMAP (McInnes et al., 2020) to reduce the dimensionality of the embeddings to a more manageable size. UMAP is a non-linear dimensionality reduction technique that preserves the local structure of the data, making it suitable for high-dimensional data such as text embeddings.
We then use HDBSCAN McInnes et al. (2017), a hierarchical density-based clustering algorithm, to group similar text data points together and identify topical structures in the data. This is particularly useful for unstructured text data, where the distribution of data points may not follow a standard distribution and noise may be present. For each cluster, we extract the five items with the highest membership probabilities, which we refer to as the cluster representatives.
Algorithm 1 Iterative Topic Taxonomy Induction
0: Document clusters $\mathcal{C}=\{C_{1},\dots,C_{n}\}$, cluster representatives $R(C_{i})$, LLM $\mathcal{M}$
0: Topic taxonomy $\mathcal{T}$ and cluster-topic assignments
1: Initialize topic list $\mathcal{T}\leftarrow\emptyset$
2:for each cluster $C_{i}\in\mathcal{C}$ do
3:  if $\mathcal{T}=\emptyset$ then
4:   Prompt $\mathcal{M}$ to generate a topic $t$ from $R(C_{i})$
5:   $\mathcal{T}\leftarrow\mathcal{T}\cup\{t\}$
6:  else
7:   Prompt $\mathcal{M}$ with $R(C_{i})$ and topics $\mathcal{T}$
8:   Constrain output to {yes, no}
9:   if no then
10:    Prompt $\mathcal{M}$ to generate a new topic $t$ from $R(C_{i})$
11:    $\mathcal{T}\leftarrow\mathcal{T}\cup\{t\}$
12:   end if
13:  end if
14:end for
15:for each cluster $C_{i}\in\mathcal{C}$ do
16:  Prompt $\mathcal{M}$ with $R(C_{i})$ and topics $\mathcal{T}$
17:  Constrain output to select exactly one topic $t\in\mathcal{T}$
18:  Assign topic $t$ to cluster $C_{i}$
19:end for
20:return $\mathcal{T}$ and cluster-topic assignments
3.2 Iterative Topic Synthesis
To generate a domain-specific label taxonomy, we use an LLM to synthesize labels from the clusters. The initial label set is empty. This list will be iteratively expanded by the LLM. If the initial list is empty, the LLM is prompted to generate a label for the first cluster.
For every other cluster, we prompt the LLM whether any of the existing labels are accurate annotations for the cluster representatives. We use constrained decoding Beurer-Kellner et al. (2024) to force the LLM to only answer with “yes" or “no". If the LLM answers “no", we prompt it to generate a new label for the cluster. We then add this label to the set of labels and repeat the process for all clusters. This iterative process continues until all clusters have been processed.
We formalize the iterative topic synthesis and annotation procedure in Algorithm 1.
3.3 Cluster Representative Labeling
The next step is to use the generated label taxonomy to annotate the cluster representatives. We prompt the LLM to generate a label for each set of cluster representatives. We once again utilize constrained decoding to limit the LLMs output to only one of the possible labels.
3.4 Supervised Classification
A downstream supervised classification task can optionally be performed using the cluster representatives and their assigned labels. This is particularly useful as clusters can be noisy, and HDBSCAN does not assign each (or even most) data points to a cluster.
For classification, we found SetFit (Tunstall et al., 2022) to be particularly effective for text classification tasks with limited labeled data. SetFit uses a two-step process: first, it fine-tunes a pre-trained sentence transformer model on the labeled data points using a contrastive loss function. Second, it trains a classification head to map the fine-tuned embeddings to the label space. This model is then used to efficiently annotate the remaining unlabeled documents.
Issue #Clusters Example Ad
economy 15 Molly Buck’s agenda is failing Iowa families. We can’t afford Molly Buck in the State House.
voting rights 13 Make Your Vote Count. RE-register to Vote in the District of Your Second Home.
crime/justice 12 Orange County Firefighters Trust Dave Min To Keep Our Communities Safe.
personal freedom 4 Tim Sheehy will always fight for Montana Values in the Senate!
voting 4 Vote Rebecca for State Rep by Nov 5th!
education 4 Vote for students. Vote for teachers. Support our public schools.
affordable housing 4 Catherine has supported more multi-family housing so we can afford homes in the communities we grew up in. Vote Stefani!
environmental protection 4 We can count on Lucy Rehm to be a responsible leader and protect our clean air and water.
immigration 3 Yadira Caraveo supported efforts to defund ICE and the border patrol. Now Colorado is being overwhelmed with illegal immigration.
election integrity 3 Question 3 will change our state’s Constitution and weaken our democracy by throwing away thousands of ballots over minor errors. Vote NO on Question 3.
abortion 2 Defend reproductive freedom - VOTE on Tuesday, November 5th!
healthcare access 1 Lucy Rehm cares about improving conditions in your local hospital. That’s why Minnesota Nurses endorsed Lucy Rehm!
property taxes 1 Caleb will work to stop reckless spending and oppose tax and fee increases.
other 2 Get the facts on the Gloucester Township sewer sale, visit VoteYesGloucesterTownship.com!
Table 1: Topics identified by the LLM in the political ads dataset. Clusters column gives the number of unique clusters assigned to each topic by the LLM.
4 Case Study
To demonstrate the effectiveness of our approach, we apply it to a large corpus of social media political advertisements from the 2024 US presidential elections. The goal is to uncover patterns the advertisers employ to target specific demographics with certain issues and moral framings.
4.1 Dataset
We collected a dataset of political advertisements running on Facebook and Instagram in the USA from October 2024. These ads were collected from the Meta Ad Library and contains $8047$ unique ads. Crucially, these ads were chosen on a day only a month away from the 2024 Presidential election, with the ad’s content covering election-related topics at both the local and federal levels. Over $99\%$ of the ads are in English, with a small fraction in Spanish, mostly in local Florida elections.
For each ad, Meta provides a unique ad ID, the ad title, ad body, and URL, ad creation time and the time span of the campaign, the Meta page authoring the ad, funding entity, and the cost of the ad (given as a range). The API also provides information on the users who have seen the ad (called ‘impressions’): the total number of impressions (given as a range and we take the average of the end points of the range), distribution over impressions broken down by gender (male, female, unknown), age ($7$ groups), and location down to the state level.
4.2 Framework Application
We apply the proposed framework to the ad texts to identify a taxonomy of political issues discussed in the ads. We embedded the ad body texts using the all-MiniLM-L6-v2 model, and removed all ads with embedding vectors too similar to another ad (cosine similarity >0.95>0.95). We then clustered these embeddings with HDBSCAN using a minimum cluster size of $15$. This resulted in $4510$ of the ads being assigned to $72$ clusters. The prompts can be found in the App.  A.
We used a locally hosted Llama-3.2-3B model Grattafiori et al. (2024) to generate the cluster labels. Experiments were conducted on a local machine using an 11th Gen Intel Core i7-11390H CPU @ 3.40GHz, without GPU acceleration. We began the topic synthesis process without any seed set, which the model expanded to $14$ total topics. For annotation, the model was additionally given the option to assign a cluster the label ‘other’ if none of the topics suited its context. Of this list, the LLM never assigned the topic ‘border security’ to any of the clusters. See Table 1 for the final list of topics and the number of clusters assigned to each topic.
Care/Harm: It suggests that someone other than the speaker is worthy of compassion or is experiencing harm, grounded in the values of kindness, tenderness, and care.
Fairness/Cheating: Emphasizes justice, personal rights, and independence; involves comparing with other groups. Advocates for equal opportunity and resists those who benefit without contributing (“Free Riders").
Loyalty/Betrayal: Based on the values of loyalty to one’s country and willingness to sacrifice for the group. Activated by a sense of unity and collective responsibility—“one for all, and all for one".
Authority/Subversion: Centers on showing respect (or resistance) toward established authority and following long-standing traditions. It includes maintaining social order and fulfilling the duties tied to hierarchical roles, such as obedience, respect, and role-based responsibilities.
Sanctity/Degradation: Beyond religion, this value highlights respect for human dignity and aversion to moral or physical corruption, promoting purity, self-control, and the belief that the body is sacred and vulnerable to defilement.
Liberty/Oppression: Captures the feelings of reactance and resentment people experience when their freedom is restricted, often leading to collective disdain for authoritarian figures and motivating unity and resistance against oppression.
Table 2: Six basic moral foundations Haidt and Graham (2007); Haidt and Joseph (2004).
4.3 Moral Foundations
In addition to topic assignment, we classify the moral foundation of each ad. Moral Foundation Theory (MFT) suggests a theoretical framework for analyzing six moral values (i.e., foundations, each with a positive and a negative polarity) central to human moral sentiment (Table 2). MFT states that political attitudes are shaped by six core moral dimensions: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, sanctity/degradation, and liberty/oppression.
To identify the moral foundation of each cluster, we first prompt the LLM to summarize the primary talking point of the cluster representatives with regard to the annotated label. The LLM is then provided with definitions of the six moral foundations, and the extracted argument is classified using constrained decoding to determine the most strongly aligned foundation. Prompts are in App. A.
4.4 Supervised Classification
To annotate the unlabeled ads, we use the labeled cluster representatives as the training set for a supervised classification task. We used a variety of classifiers, including Logistic Regression, XGBoost Chen and Guestrin (2016), RoBERTa Liu et al. (2019), and SetFit Tunstall et al. (2022).
For the labeled ads, we took all ads which HDBSCAN assigned a high membership probability to (greater than $0.98$), which resulted in a total of $2337$ ads, of which the four majority classes ‘voting rights’ (26%), ‘crime/justice’ (19%), ‘education’ (17%), and ‘abortion’ (11%) made up 3 quarters (73%) of the labeled ads. The remaining classes were much smaller, with the smallest class being ‘healthcare access’ and ‘property taxes’ (0.64% each).
We randomly selected $200$ ads from the unassigned clusters, and annotated them manually. We used these annotations as the ground truth labels for the classification task.
For each model, we used the all-MiniLM-L6-v2 model to embed the ad texts and employed GridSearchCV to optimize the hyperparameters. We used $5$-fold cross-validation to evaluate the models, and selected the best-performing model based on the macro F1 score. The results of this evaluation are shown in Table 3.
Model Macro F1 Accuracy
Logistic Regression 0.37 0.55
XGBoost 0.31 0.53
RoBERTa 0.32 0.53
SetFit 0.36 0.6
Table 3: Classification performance across annotated ads for topic assignment (15 unbalanced classes).
4.5 Demographic Targeting Analysis
Using the ad impressions data, we can analyze the targeting data of the ads. The analysis focuses on the use of social media platforms, particularly Meta and Instagram, to disseminate targeted political advertisements. We examine the strategies employed by the advertisers to reach specific demographic groups. Meta allows advertisers to define their audiences on the basis of location (state-level), age and gender. We utilized this data, together with the annotations, to analyze the microtargeting strategies of the political entities.
To understand how messages are tailored to different demographics, we look at the positive pointwise mutual information (PPMI) Church and Hanks (1990) between the words in the advertisements and the demographic groups. The PPMI is a measure of association between two events, in this case, the words in the advertisements and the demographic groups. A higher PPMI value indicates a stronger association between a word and a demographic group. We compare both the topics discussed in the advertisements and the moral foundations that are used to appeal to the different demographic groups. This annotated corpus enables detailed analysis of issue emphasis and demographic targeting, reported in Section 5.
5 Results and Analysis
We evaluate the performance of our approach on the social media advertisement dataset, evaluating the performance of both annotation and classification tasks. Additionally, we discuss the results of our case study, showcasing differences in moral framing and issue distribution across different demographics. Our analysis proceeds at two complementary levels: (i) cluster-level analysis, which characterizes issue salience and moral framing using LLM-synthesized topic clusters, and (ii) ad-level analysis, which leverages individual ad annotations to examine spending patterns, funding style, and moral framing at scale.
Model Average Score Best Label
BERTopic 1.2 3
TopicGPT-style 2.7 5
Our Method 2.8 12
Table 4: Annotation results. The average score is out of 5, and the best label is the number of times that model’s label was selected as the most fitting (among all annotators).
(a) Proposed Framework Topics
(b) LDA Topics
Figure 2: Correlations between moral foundations and topics. Color intensity indicates the strength of the correlation.
5.1 Baseline Comparison
First, we evaluate the topic labels by comparing the performance of our proposed framework with the baseline BERTopic  Grootendorst (2022). For both models, the same clusters are used, and both models use the same LLM to generate the topic labels. For BERTopic, the LLM generated labels based on both the representative documents and the top 10 keywords.
To assess whether iterative taxonomy construction provides benefits beyond direct LLM labeling, we introduce an additional baseline inspired by TopicGPT Pham et al. (2023). For each cluster, we prompt the same LLM with the cluster representative ads and ask it to generate a concise topic label in a single step, without access to previously generated labels or iterative refinement. This baseline isolates the effect of structured taxonomy growth and constrained validation in our framework.
Two annotators were each given $10$ randomly selected clustered ads. The annotators were volunteer graduate students and faculty who spoke fluent English with backgrounds in computational social science, natural language processing, and computer science. For each ad, they were given both the label from BERTopic and our proposed framework (the order of the labels was randomized). The annotators were asked to select the label that best represented the ad, and to provide a score from 1-5 for each label, where $5$ would be a perfect fit, and $1$ would describe a totally irrelevant label. The results of this comparison are shown in Table 4.
These results suggest that the proposed framework yields more consistent and better-aligned topic labels than both BERTopic and the single-shot TopicGPT-style baseline under this evaluation setup. A Cohen’s Kappa Cohen (1960) score of $0.66$ was achieved between the two annotators, indicating a moderate level of agreement. For $5$ out of the $20$ rounds, the annotators gave the same score for both models (in all cases a score of 1).
Analyses.
The following analyses are intended to illustrate the types of downstream insights enabled by the induced taxonomy, rather than to make exhaustive claims about political advertising behavior.
5.2 Topic Salience and Moral Framing
We present some selected results from our case study. Firstly, we analyze the correlation between moral foundations and the latent topics, taking LDA topics as a baseline. To compute the correlation, we first one-hot encoded each ad twice: once by topic and once by moral foundation. We then computed the pairwise correlation using the Pearson correlation between these one-hot encodings. The results are shown in Fig. 2. In Fig. 2(a), we can see very strong correlations between certain moral foundations and topics. For example, the Fairness/Cheating moral axis is strongly correlated with the crime/justice topic, while the Loyalty/Betrayal moral axis is strongly correlated with the immigration topic. This suggests that the induced topics align with moral framing patterns in this dataset. In contrast, the LDA topics in Fig. 2(b) show much weaker correlations with the moral foundations.
We analyze demographic targeting using positive PPMI between ads and audience attributes (Fig. 4 in App. B). Clear age- and location-specific patterns emerge: in Florida, younger audiences are disproportionately shown affordable housing ads, while older audiences receive more abortion-focused messaging. Geographic contrasts are also evident: male audiences in Montana are more exposed to environmental protection and personal freedom ads, whereas male audiences in Virginia are targeted more heavily with crime/justice and voting-related content, reflecting systematic alignment between issue framing and demographic context.
Figure 3: Average spend and impressions reveal which topics dominate the political ad landscape.
5.3 Ad-Level Analysis
To enable finer-grained analysis beyond cluster-level structure, we annotate individual ad using the synthesized topic taxonomy. Using the generated topics as labels, we apply an LLM-based annotation (model: mistralai/Mistral-7B-v0.1 Jiang et al. (2023)) to assign each ad a topic and moral foundation. This enables ad-level analysis of spending, exposure, and funder behavior, which we report below.
5.3.1 Spending and Reach by Topic
We begin with a descriptive analysis of ad spending and exposure across topics. Fig. 3 shows the average spend (left panel) and average impressions (right panel) per topic. Topics such as abortion, personal freedom, and voting rights receive the highest average spend, indicating strong advertiser investment in normatively salient and mobilizing issues. In contrast, immigration and border security achieve comparatively high impressions with lower average spend, suggesting broader organic reach or more cost-efficient targeting. Issues such as property taxes and local services (e.g., sewer sale) receive minimal investment and exposure. Overall, the divergence between spend and impressions highlights strategic differences in how advertisers allocate resources versus how effectively topics scale to large audiences.
5.3.2 Top Funders by Issue Domain
We analyze the concentration of advertising spend across funding entities for three central issues of 2024 US presidential election—economy, abortion, and crime/justice [^2] https://www.pewresearch.org/politics/2024/09/09/issues-and-the-2024-election/ (Fig. 5, App. C.1).
Spending on the economy is highly concentrated, with one dominant national organization and a sharp drop-off among other major PACs. Abortion exhibits the strongest concentration overall, with a small number of opposing organizations accounting for the majority of spending, underscoring its role as a high-salience wedge issue. In contrast, crime/justice shows a more distributed funding profile, with multiple PACs contributing comparable levels of investment, often tied to specific states or jurisdictions. App. C.1 provides the details. Overall, these patterns reveal substantial variation in funding concentration across issues: economy and abortion messaging are driven by a small set of major actors, whereas crime/justice advertising reflects more decentralized and localized campaign strategies.
5.3.3 Moral Framing and Ad Spending
We examine how moral framing aligns with financial investment across issue domains. Aggregate spending by moral foundation (Fig. 6, App. C.2) reveals distinct moral structures across topics. Economy advertising distributes spending across multiple foundations—primarily Fairness/Cheating and Liberty/Oppression—indicating moral plurality. In contrast, abortion advertising concentrates spending on two sharply opposed foundations, Liberty/Oppression and Sanctity/Degradation, reflecting moral polarization.
Funder-level breakdowns (Fig. 7, App. C.2) show that these aggregate patterns arise from strategic specialization by major funding entities rather than uniform framing across actors. Detail in App. C.2
6 Conclusion
We propose a general, seed-free framework for iterative topic taxonomy induction using LLMs. By combining embedding-based clustering with constrained, multi-pass LLM inference, the framework enforces global consistency across clusters and improves topic label quality relative to both classical topic models and single-shot LLM labeling under human evaluation. We validate the framework through a large-scale political advertising case study, demonstrating how the induced taxonomies enable interpretable downstream analyses such as issue prevalence, moral framing, and demographic targeting in this setting. While our empirical evaluation focuses on electoral advertising, the framework itself is not domain-specific and can be applied to other large, unlabeled text corpora. An important direction for future work is to evaluate taxonomy stability across domains, clustering granularities, and LLM configurations, and to further study the robustness of iterative synthesis under different prompting and ordering conditions.
7 Limitations
While our framework reduces manual annotation effort and generalizes across domains, several limitations remain. Firstly, we purposefully choose a smaller LLM to show that our framework is not limited to computational resources or budget constraints. However, the performance of our framework is likely to improve with more powerful LLMs.
Secondly, while we demonstrate the effectiveness of our framework on a political advertisement corpus, it may not generalize to all domains. In particular, domains with highly overlapping or abstract topic boundaries may require additional constraints or hierarchical structure.
Our framework relies on LLMs, which may introduce biases or hallucinate misleading topic labels. Additionally, automatic interpretation of political content may risk oversimplification or misrepresentation of nuanced discourse.
Future work should explore the applicability of our framework across different domains and tasks. Clustering performance can be sensitive to embedding quality and density-based parameters, and we recommend further investigation into the optimal parameters for different datasets.
8 Ethical Consideration
To the best of our knowledge, we did not violate any ethical code while conducting the research work described in this paper. We report the technical details for the reproducibility of the results. The author’s personal views are not represented in any results we report, as it is solely outcomes derived from machine learning and/or AI models. The data collected in this work was made publicly available by the Meta Ad Library API. The data do not contain personally identifiable information and report engagement patterns at an aggregate level.
Appendix A Prompts
Below are the prompts used in the various steps of the framework. Constrained decoding was implemented with the guidance framework [^3] https://github.com/guidance-ai/guidance.
System Prompt
You are a political analyst. You are given the following set of political ads to analyze. You will assign the ads a topic that best summarizes the key issue discussed in these ads.
A.1 Topic Synthesis
Binary Output
Is there a topic in the following list that well describes the key issue discussed in these ads?
Topics:
{{ topics }}
Can the ads can be summarized by one of the previous topics: Answer with "yes" or "no".
Generate New Topic (User)
In three words or less, describe the issue that ads discuss, summarizing the topic. Examples: "Abortion", "Climate Change", etc.
Output constriained to ["yes", "no"].
Generate New Topic (Assistant)
A better topic for these ads is: "
The LLM would finish the above, with the quote (") as a stop token.
A.2 Annotation
System Prompt
You are a political analyst. You are given the following set of political ads to analyze: {{ ads }}
Select Topic (User)
Summarize the main talking point of the ads. Do so in a single sentence.
Output constrained to set of topics.
Summarize (User)
Summarize the main talking point of the ads. Do so in a single sentence.
Select Moral Foundation (User)
Which of Which of the following moral foundations best describes the arguments used in the ads about {{ topic }}?’
{{ moral_foundations_with_definitions }}
(a) Females: AL vs WA
(b) All Age groups: FL
(c) Males: MT vs VA
(d) Moral Framing across States
Figure 4: Heatmaps showing the PPMI analysis of moral framings and political issues across different demographics. The color intensity indicates the strength of the correlation.
Appendix B Substantive Insights: Topic Salience and Moral Framing
We analyze the distribution of moral framings and political issues across different demographics to uncover political microtargeting strategies. We used the positive point-wise mutual information (PPMI) to detect topics that appear disproportionately in ads shown to specific demographics, potentially indicating targeted messaging. We visualize the results of the PPMI analysis in Figure 4. The heatmaps show the distribution of moral framings and political issues across different demographics. Some interesting observations include the differences in ads targeting different states, or the shift in topics among different age groups in Florida (Figure 4(b)). For example, young people in Florida are more likely to be targeted with ads related to affordable housing, while older people are more likely to be targeted with ads related to abortion. This shows generational targeting strategies, where economic concerns are emphasized for younger groups, while older groups are mobilized on moralized issues like abortion.
On the other hand, Figure 4(c) shows that environmental protection and personal freedom are heavily overrepresented among male audiences in Montana, while abortion also appears moderately elevated. By contrast, male audiences in Virginia are disproportionately exposed to crime/justice and voting-related advertising. These contrasts highlight how campaign strategies selectively target localized demographic niches: progressive-leaning issue frames resonate more strongly in Montana male audiences, whereas mobilization around law-and-order and electoral process is more pronounced among Virginia males.
Appendix C Substantive Insights: Ad-Level Analysis
Using the synthesized topics, we annotate individual ads to enable ad-level analysis. Cluster-level annotations reveal latent issue and moral structure, while ad-level annotations enable analysis of spending, funder concentration, and targeting strategies.
Figure 5: Top five funding entities by total advertising spend for three major issue areas (Economy, Abortion, and Crime/Justice). The figure shows substantial concentration of spending within each topic, with one or two organizations accounting for a large share of total investment, particularly for abortion and economic messaging, while crime/justice exhibits a comparatively more distributed funding profile across multiple PACs.
C.1 Top Funders
We examine the concentration of advertising spend among funding entities across major issue domains. Fig. 5 presents the top $5$ funding entities by total spend for three central issues in the 2024 election—economy, abortion, and crime/justice [^4] https://www.pewresearch.org/politics/2024/09/09/issues-and-the-2024-election/.
For the economy, spending is highly concentrated among a small number of well-funded national organizations. One entity dominates overall investment, with a sharp drop-off among the remaining top funders. Other contributors—including large advocacy and leadership PACs—account for substantially smaller but still nontrivial shares. This pattern suggests that economic messaging is driven primarily by a limited set of organizations with the capacity to sustain large-scale, sustained advertising campaigns.
The abortion domain exhibits the strongest concentration of spending across the three topics. A single organization accounts for a disproportionate share of total expenditure, while the remaining top funders contribute at comparatively similar but much lower levels. The presence of multiple organizations aligned with opposing positions indicates intense contestation, but the highly skewed spending distribution highlights the outsized role of a small number of actors in shaping abortion-related advertising at scale. This reinforces abortion’s status as a high-salience wedge issue in paid political communication. For example: Florida Voters Against Extremism [^5] https://dos.elections.myflorida.com/committees/ComDetail.asp?account=84315 is the largest spender, anchoring anti-abortion rights mobilization, particularly around state-level ballot initiatives. In contrast, progressive-aligned group such as Montanans Securing Reproductive Rights [^6] https://www.influencewatch.org/organization/montanans-securing-reproductive-rights/ invests heavily in support of abortion rights.
In contrast, the crime and justice topic displays a more distributed funding profile. While two PACs emerge as the largest spenders, the remaining top entities contribute comparable levels of investment, resulting in a less sharply skewed distribution. Several of these organizations are tied to specific states or jurisdictions, indicating a mix of national framing and localized campaign investment. This pattern suggests that crime and justice advertising is both financially intensive and geographically fragmented, reflecting its relevance to state- and local-level electoral contexts.
Overall, Fig. 5 illustrates substantial variation in funding concentration across issue domains. While economy and especially abortion advertising are dominated by a small number of major organizations, crime/justice messaging involves a broader set of actors with more evenly distributed spending, pointing to distinct strategic and organizational dynamics across policy areas.
Figure 6: Aggregate ad spend by moral foundation shows distinct moral structures across topics: abortion centers on liberty/oppression and sanctity/degradation, crime/justice on loyalty/betrayal and authority/subversion, and the economy on fairness/cheating and liberty/oppression.
C.2 Moral Foundations in Funded Messaging
We link moral framing to financial investment across issue domains. Fig. 6 presents total advertising spend by moral foundation for three high-salience topics: abortion, crime/justice, and economy. This aggregate view reveals clear differences in moral structure across issue domains, independent of individual funding entities.
We next examine how these aggregate moral patterns emerge from individual funding entities. Fig. 7 compares the top $5$ funding entities in two salient topics—economy and abortion disaggregated by dominant moral foundation appeals. For the economy, spending is distributed across multiple moral foundations, with substantial investment in both Fairness/Cheating and Liberty/Oppression frames. Several major funders emphasize fairness-oriented rhetoric (e.g., taxation, economic equity), while others concentrate spending almost exclusively on liberty-based appeals, particularly around regulation, government intervention, and market freedom. Minor contributions to Loyalty/Betrayal framing appear for some organizations but remain secondary. Overall, economic messaging exhibits clear moral plurality, with different actors strategically emphasizing distinct moral narratives rather than converging on a single frame.
In contrast, the abortion domain displays a markedly different moral profile. Spending is dominated by two moral foundations—Liberty/Oppression and Sanctity/Degradation with funders sharply specializing in one or the other. Organizations aligned with reproductive rights primarily invest in liberty-based appeals centered on bodily autonomy and government overreach, whereas anti-abortion organizations concentrate heavily on sanctity-based framing. Contributions invoking Care/Harm or Loyalty/Betrayal are present but comparatively limited. This pattern indicates that abortion advertising is characterized not by moral diversity within actors, but by moral polarization across actors.
Taken together, Fig.  7 shows that moral framing is closely tied to both issue domain and funder strategy: economic advertising supports multiple coexisting moral narratives, while abortion advertising is structured around a small number of sharply opposed moral foundations.
Figure 7: Funder-level moral framing for economy and abortion ads: economic messaging spans multiple moral foundations across funders, while abortion advertising is dominated by sharply polarized Liberty/Oppression and Sanctity/Degradation frames.